OpenStudy (anonymous):

evaluate lim : cos theta/sin 2theta theta-pi/2

OpenStudy (ujjwal):

sin2theta=2sintheta*costheta

OpenStudy (ujjwal):

can you do further?

OpenStudy (anonymous):

no i cant actually but is wat i get csc(2theta)cos(theta)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 90 }\cos \theta /\sin2 \theta \] we know that \[\sin 2\theta=2 \sin \theta \cos \theta\] \[\lim_{\theta \rightarrow 90 } \cos \theta / 2 \sin \theta \cos \theta\] \[\lim_{ \theta \rightarrow 90} 1/ 2 \sin \theta\] now apply limt\[1 / 2\sin 90\] where sin90=1 so 1/2(1) 1/2 ans

OpenStudy (anonymous):

Now any problem

OpenStudy (anonymous):